• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name riordan

    Log entry time 07:56:15 on March 05, 2011

    Entry number 350250

    keyword=Owl Shift Summary

    SL:  Seamus Riordan
    TO:  Megan Friend
    3rd: David Anez
    
    00:00   Start shift - Setting #6 Al dummy in progress
    00:58   Ask for escorted access to move LHRS
    01:35   LHRS 14 degree pin in place - moving to 46 degrees
    02:01   LHRS at 46 degrees, starting at 40uA on 4cm LH2
    02:09   Turned on FFB
    02:50   Start production on 4cm LH2 at 80uA
    04:00   Move to 4cm Al dummy
    04:05   Start 1hr Al
    05:10   BeO pointing run
    05:20   Back to LH2 production
    
    
    Good Production/Calibration Runs
    Setting #6
    2360    20uA 4cm Al dummy        800k events
    Setting #7
    2366    80uA 4cm LH2               2M events
    2367    80uA 4cm LH2               2M events
    2368    80uA 4cm LH2               2M events
    2369    80uA 4cm LH2               2M events
    2370    20uA 4cm Al dummy        713k events
    2371    20uA BeO                  60k events
    2372    80uA 4cm LH2               2M events
    2373    80uA 4cm LH2               2M events
    2374    80uA 4cm LH2               2M events
    2375    80uA 4cm LH2               2M events
    2376    80uA 4cm LH2               2M events
    2377    80uA 4cm LH2               2M events
    2378    80uA 4cm LH2               2M events
    2379    80uA 4cm LH2               In progress
    
    Total charge for setting #6:       1.24 C
    Accumulated charge for setting #7: 0.97 C